I am not ggod at word problems and I am having trouble on this one

The eqaution H=-16t^2+112t gives the height of an arrow, shot upward from the ground with initial velocity of 112 ft/s, where t is the time after the arrow leaves the ground. find the time it takes for the arrow to reach a height of 180ft.

Thanks a million

Answer by stanbon(75887)

The equation H=-16t^2+112t gives the height of an arrow, shot upward from the ground with initial velocity of 112 ft/s, where t is the time after the arrow leaves the ground. find the time it takes for the arrow to reach a height of 180ft.
--------------
Let H=180 and solve for t:
-16t^2+112t-180=0
t^2-7t+12 = 0
(t-3)(t-4)=0
t =3 seconds or t = 4 seconds
The 3 seconds is on the way up; the 4 seconds is on the way down.
================
Cheers,
Stan H.